Which of the following pentapeptides will exhibit the highest extinction coefficient at 280 nm? Assume that each tryptophan has an extinction coefficient of 5690 M-1cm-1, tyrosine has an extinction coefficient of 1280 M-1cm-1, and the absorbance of phenylalanine at 280 nm is negligible.

TEDSE
#TYWSF
RYYFE
FEWFE

The UV-vis absorption spectrum of the peptide RYYFE collected in a 1 cm pathlength cuvette exhibits an absorbance of 0.050 at 280 nm. What is the concentration (in micromolar) of the peptide in this sample?
Answer #1

W is tryptophan and Y is tyrosine.

Tryptophan and tyrosine has extinction coefficient of 5690 and 1280 M-1cm-1.

The pentapeptide with larger number of this two will have greater extinction coefficient.

TYWSF has 1 tryptophan and 1 tyrosine.while RYYFE has two tyrosine and FEWFE has 1 tryptophan.

Thus,TYWSF has highest extinction coefficient due to both tryptophan and tyrosine.

RYYFE has two tyrosine, so the extinction coefficient is 1290×2=2580 M-1cm-1.

We know, A=€cl

0.05=2580×c×1.

c=0.05/2580

c=1.94 E-5 M

c=19.4 micromolar.